A sum of money at simple interest amount Rs.815 in 3 years and Rs.854 in 4 years what is the principle amount invested?
Answer by Theo(13342)

you have 815 in 3 years and 854 in 4 years.
thie means the increase is equal to 39 in 1 year
3 * 39 = 117
815 - 117 = 698
your principal is equal to 698.
in 4 years the money grows at the same amount each year.
time point 0 = 698
time point 1 = 737
time point 2 = 776
time point 3 = 815
time point 4 = 854
your simple interest rate is equal to 39 / 698 = .0558739255.
the simple interest is calculated off of the original principal and the same amount of interest is added each year.